OpenStudy (anonymous):

A mixture of 30 pounds of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of the $1.50 chocolate were used to make the mixture?

OpenStudy (anonymous):

Price Equation: 1.5*a + 0.9*b = 1.1*30 = 33 Quantity Equation: a + b = 30 Solve the system.

OpenStudy (anonymous):

a=10 b=20

OpenStudy (anonymous):

choices are 15 10 12